# Gravitational waves from vacuum bubbles: ultraviolet dependence on wall thickness

### v1

@@ Abstract
The gravitational wave (GW) spectrum from the first-order phase transition can be characterized by a few phenomenological parameters but with high degeneracies in model/data distinguishments. In this paper, we look into the high-frequency power law of the GW spectrum with preliminary numerical simulations for both quantum and semi-classical pictures of vacuum decay. We first reveal an anti-correlation of the high-frequency power law to a certain power of the ratio between the wall thickness and bubble radius at the onset of bubble collisions, which can be further approximated analytically by some other phenomenological model characteristics to break the model degeneracy.
@@ I Introduction
As the early Universe evolved, the spontaneous breaking of a continuous symmetry may have led to a cosmological first-order phase transition (FOPT) Mazumdar and White (2019); Hindmarsh et al. (2021). At the beginning of an FOPT, cosmic bubbles separating false and true vacua can nucleate stochastically via vacuum decay in field theory. In addition to the well-known channel via quantum tunneling Steinhardt (1982); Coleman (1977); Callan and Coleman (1977), a semi-classical one is also plausible to realize vacuum decay Braden et al. (2019). The former is purely a quantum effect allowing for barrier penetration with lower energy, while the latter could originate from the classical evolution of occasionally developed perturbations Hertzberg and Yamada (2019); Blanco-Pillado et al. (2019); Wang (2019); Huang and Ford (2022). After nucleations, bubbles expand rapidly Cai and Wang (2021); Wang and Yuwen (2023); Wang et al. (2024a); Yuwen et al. (2024) and finally percolate via violent collisions, during which stochastic gravitational wave backgrounds (SGWBs) are generated Caprini et al. (2016, 2020) via bubble collisions Witten (1984); Hogan (1986), sound waves Hogan (1986), and magneto-hydrodynamic (MHD) turbulences Witten (1984) to be detectable by currently running GW detections (LIGO-Virgo Romero et al. (2021); Huang et al. (2021); Jiang and Huang (2023); Badger et al. (2023); Yu and Wang (2023), PTAs Xu et al. (2023); Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023)) and future planning GW detectors (LISA Amaro-Seoane et al. (2017); Baker et al. (2019), Taiji Hu and Wu (2017); Ruan et al. (2020), TianQin Mei et al. (2021); Luo et al. (2016)).
@@
However, the high-frequency tail in the GW power spectrum has not been fully understood yet. Taking the simplest case of vacuum phase transitions without thermal fluids as an example, previous analytical modelings of thin-wall collisions with Jinno and Takimoto (2017) and without Jinno and Takimoto (2019) the envelope approximation agree on an $k^{-1}$ power law for the high-frequency part of the GW spectrum, but a semi-analytical simulation from a dubbed bulk flow model Konstandin (2018) also beyond the envelope approximation produces rather different $k^{-2}$ and $k^{-3}$ power laws at ultraviolet (UV) for ultrarelativistic and nonrelativistic walls, respectively. Further relaxing the thin-wall approximation, a previous numerical simulation Cutting et al. (2021) has found a steeper UV power than $k^{-1}$ depending roughly on a model parameter that is intimately related to the bubble wall thickness.
In this paper, we will make this relation clearer by preliminarily quantifying this relation between the high-frequency power law of GW spectrum and bubble wall thickness at collision. Following Refs. Braden et al. (2015); Jinno et al. (2019); Lewicki and Vaskonen (2020); Cutting et al. (2021), we first numerically simulate the bubble nucleations, expansion, and collisions in Sec. II and Sec. IV, and then calculate and extract the GW spectrum in Sec. V. In doing so, we analytically derive in Sec. III for the first time the mean bubble separation at nucleation for runaway bubbles from vacuum phase transitions, where bubbles continue to accelerate without approaching a constant terminal velocity at collisions. Regardless of the nucleation channels of quantum or semi-classical types, we find in Sec. VI the high-frequency power law decreases to a certain power of the wall thickness at collisions. Finally in Sec. VII, based on the total energy conservation, we analytically approximate the wall thickness at collisions as a function of other more model-independent parameters, such as the phase hierarchy $\Delta\phi/M_{\mathrm{Pl}}$, strength factor $\alpha=\Delta V/\rho_{\mathrm{rad}}$, inverse duration $\beta/H$, and wall velocity $v_{w}$ as we will define later. This informative modeling of the high-frequency power law together with the remaining parts of the GW spectrum largely breaks model degeneracies for future GW constraints on FOPTs. The last section VIII is devoted to conclusions and discussions.
@@ II Bubble nucleation
In addition to quantum tunneling, bubble nucleation can also occur through a semiclassical mechanism Braden et al. (2019); Hertzberg and Yamada (2019); Blanco-Pillado et al. (2019); Wang (2019); Huang and Ford (2022). In this scenario, quantum fluctuations Hertzberg and Yamada (2019), thermal fluctuations Berera and Fang (1995) or gravitational effects Blanco-Pillado et al. (2019) drive the scalar field or its time derivative to develop a non-zero profile. The classical evolution of these fluctuations then leads to the bubble formation in real space instead of the Euclidean space Braden et al. (2019). One of the simplest realizations Blanco-Pillado et al. (2019); Wang (2019) of this semiclassical mechanism is the time derivative of the scalar field occasionally develops a Gaussian profile large enough to eventually overcome the potential barrier directly from a homogeneous field at the false vacuum. The initial condition for bubble formation is given by
@@
where $A_{0}=\sqrt{2eV_{b}}$ is defined in such a way that the critical velocity $\dot{\phi}_{c}\equiv\sqrt{2V_{b}}$ is realized exactly at $r=R$ for $A=A_{0}$. The profiles of (6) are shown in the right panel of Fig. (2). Ref. Wang (2019) provides a detailed discussion on the ranges of $A/A_{0}$ and $R/R_{0}$ that permit bubble nucleation. In our subsequent calculations, we exactly follow the same configuration and procedure as Ref. Wang (2019) to generate the semi-classical bubble from simply fixing $A/A_{0}=1$ and $R/R_{0}=1$ within the range where bubble nucleation can occur. The specific choice of parameters $A$ and $R$ within this allowable range does not affect our final results.
@@ III Bubble separation
After obtaining the scalar field initial profiles from quantum tunneling and semiclassical mechanisms, there is one last quantity to be determined for the initial configuration of numerical simulation: the mean bubble separation at the nucleation time, which should be twice the mean bubble radius at collisions. In this study, we focus on the vacuum FOPT, and hence, the potential (1) is independent of the background temperature. In such a scenario, the bubble nucleation rate $\Gamma$ is time-independent and given by Coleman (1977); Callan and Coleman (1977); Linde (1981, 1983); Cai et al. (2017b)
$\Gamma=\frac{1}{R_{0}^{4}}\left(\frac{S_{E}}{2\pi}\right)^{2}e^{-S_{E}},$ (7)
@@
In the vacuum FOPT, the bubble expands without bound and runs away to approach the speed of light, and hence, there is no constant terminal wall velocity $v_{w}$ by the time of bubble collisions. In this case, the false vacumm fraction $F(t)$ is expressed as Guth and Weinberg (1983); Turner et al. (1992)
@@
where $\Gamma_{\rm E}$ is the Euler-Gamma function and $\Gamma_{\rm E}(8/7)\simeq 0.935$. In Table. 1, we present the mean bubble separation for various values of the potential parameter $\omega$. It is observed that, when $\omega<0.6$, the bubble separation becomes extremely large, and simulating such a system would require significant computational resources. Therefore, in this study, we restrict our simulations to values of $\omega$ ranging from 0.6 to 1.0, with an interval of 0.05.
@@
Table 1: The mean bubble distance $d$ at the nucleation time in the quantum phase transtion for different $\omega$ values.
Lastly, our analysis above focuses on the mean bubble separation in the quantum phase transition, which is determined by the bubble nucleation rate. According to Ref. Braden et al. (2019), the nucleation rate in the semiclassical phase transition is nearly identical to that in the quantum phase transition. Therefore, when simulating the semiclassical phase transition, we will also use the same bubble separation values presented in Tab. 1.
@@ IV Bubble expansion and collision
but now both evolve in the real space with initial profiles containing two bubbles. For simplicity, we assume the two bubbles are identical. When the bubbles are separated by a distance $d$, their centers can always be positioned on the $z$-axis at $z=\pm d/2$ by adjusting the coordinate system. Now, the evolution of this two-bubble system is best described in cylindrical coordinates $(t,\rho,\theta,z)$, in which case the equation (20) is expanded as
@@
During our simulation, we discretize the space using a lattice with 1600 grid points per spatial dimension. The entire range for $z$ and $\rho$ is set between $-1.5d$ and $1.5d$, and from 0 to $1.5d$, respectively. For the time dimension, we divide it into $10^{5}$ slices and simulate the bubble evolution using the sixth-order Runge-Kutta method.
@@
where $\mathrm{d}\mathcal{V}$ is the volume element integrated over the open regions of true bubbles. Throughout the numerical simulations, the total energy $E_{T}\equiv E_{K}+E_{G}+E_{V}$ should remain conserved (or vanished specifically for the quantum phase transition). One of our numerical results for the time evolution of semiclassical bubble expansion is shown in Fig. 3, where it can be seen that $E_{K}$, $E_{G}$, and the absolute value of $E_{V}$ are converged into a constantly evolving total energy $E_{T}$. For quantum phase transitions, the numerical stability has also been checked with $E_{T}=0$ since the converted potential energy from the false vacuum to true vacuum has all been deposited into the kinetic and gradient energies of the bubble wall.
@@
After the bubble collision, the overlapping region of the two bubbles can rebound back and forth the potential barrier, similar to the trapping effect discussed in Ref. Jinno et al. (2019). When the bubble walls are thick (middle row of Fig. 4), the trapping effect is weakened, but oscillations around the true vacuum become pronounced since the exit point is now on the potential hill away from the true vacuum, generating outgoing waves that follow the expanding bubble walls. As the wall thickness further increases (bottom row of Fig. 4), a significant difference emerges in the semiclassical phase transition, where the scalar field evolution seems to resemble the expansion of a single bubble after collision. This phenomenon occurs as a large wall thickness corresponds to a small bubble separation (as shown in Tab. 1), causing the classical perturbations to be so close that they can be nearly treated as a single perturbation. In this case, gravitational waves should not be generated, but it is still intriguing for future study as a new way to generate the bubble, not from nucleation, but from the collision of false-vacuum classical perturbations. Nevertheless, in the following Sec. V and Sec. VI, we consider the semiclassical phase transition only for the case where two bubbles remain distinct (corresponding to $\omega<0.9$).
@@ V Gravitational waves
In this work, we consider the system of two identical bubbles, and the cylindrical symmetry in this system allows the integral over the azimuthal angle to be performed analytically. Furthermore, without loss of generality, we can set $\hat{\mathbf{k}}=(\sin\theta_{k},0,\cos\theta_{k})$. The non-zero components of $T_{ij}$ are expressed as
$T_{\rho\rho}(\theta_{k},k)$ $=-\frac{1}{2}\int\mathrm{d}te^{ikt}\int_{0}^{\infty}\mathrm{d}z\cos\left(k\cos\theta_{k}z\right)\int_{0}^{\infty}\mathrm{d}\rho\ \rho\left[\sin^{2}\theta_{k}J_{0}\left(k\sin\theta_{k}\rho\right)+\left(\cos^{2}\theta_{k}+1\right)J_{2}\left(k\sin\theta_{k}\rho\right)\right](\partial_{\rho}\phi)^{2},$
$T_{\rho z}(\theta_{k},k)$ $=-\int\mathrm{d}te^{ikt}\int_{0}^{\infty}\mathrm{d}z\sin\left(k\cos\theta_{k}z\right)\int_{0}^{\infty}\mathrm{d}\rho\ \rho J_{1}\left(k\sin\theta_{k}\rho\right)\partial_{\rho}\phi\partial_{z}\phi,$
$T_{zz}(\theta_{k},k)$ $=\int\mathrm{d}te^{ikt}\int_{0}^{\infty}\mathrm{d}z\cos\left(k\cos\theta_{k}z\right)\int_{0}^{\infty}\mathrm{d}\rho\ \rho J_{0}\left(k\sin\theta_{k}\rho\right)(\partial_{z}\phi)^{2},$ (28)
@@
$\omega$ $k_{c}d/2$ $E_{\rm p}$ $A_{m}/A_{n}$ $a_{m}$ $a_{n}$ $l_{\rm col}/R_{\rm col}$ 0.60 3.27 1.00 3.32 2.94 1.73 0.0084 0.65 2.72 0.52 3.12 2.94 1.45 0.0151 0.70 2.36 0.32 3.12 2.93 1.28 0.0254 0.75 2.26 0.18 3.21 2.93 1.19 0.0346 0.80 2.15 0.12 3.32 2.93 1.10 0.0455 0.85 1.96 0.11 3.23 2.92 1.01 0.0546
Table 2: The numerical fitting results for the quantum phase transition and semiclassical phase transition. The first column is the potential parameter $\omega$. Parameters from the 2nd to the 6th columns are those shown in the parameterization (30) of the GW spectrum. Here, we set $E_{\rm p}$ in the semiclassical phase transition with $\omega=0.6$ as 1 to normalize the peak value of GW spectra. The 7th column is the ratio of the bubble thickness to its radius at the collision time.
@@
where $k_{c}$ denotes the peak frequency at which the spectrum reaches its maximum $E_{\rm p}$. Here, $A_{m}$, $A_{n}$, $a_{m}$, and $a_{n}$ characterize the power-law behavior of the GW spectrum. When $k\ll k_{c}$, the spectrum $\Omega_{\rm GW}$ follows the power law $k\mathrm{d}E/\mathrm{d}k\sim(k/k_{c})^{a_{m}}$, whereas for $k\gg k_{c}$, it follows $k\mathrm{d}E/\mathrm{d}k\sim(k/k_{c})^{-a_{n}}$. Numerical fitting results for the parameters in (30) are presented in Table. 2. It is easy to see that, at low frequencies, the $a_{m}$ values for different $\omega$ in both phase transitions are nearly identical, $a_{m}\approx 3$, as expected from the causality requirement Cai et al. (2020). At high frequencies, however, the $a_{n}$ values vary depending on $\omega$ and the type of phase transition. Additionally, the peak values $E_{\rm p}$ of the spectrum for both phase transitions decreases as the bubble wall thickness increases, corresponding to higher $\omega$ values.

### v3

@@ Abstract
The gravitational wave (GW) spectrum from the first-order phase transition can be characterized by a few phenomenological parameters but with high degeneracies in model/data distinguishments. In this paper, we look into the high-frequency power law of the GW spectrum with preliminary numerical simulations for both quantum and semiclassical pictures of vacuum decay. We first reveal an anticorrelation of the high-frequency power law to a certain power of the ratio between the wall thickness and bubble radius at the onset of bubble collisions, which can be further approximated analytically by some other phenomenological model characteristics to break the model degeneracy.
@@ I Introduction
As the early Universe evolved, the spontaneous breaking of a continuous symmetry may have led to a cosmological first-order phase transition (FOPT) Mazumdar and White (2019); Hindmarsh et al. (2021); Athron et al. (2024). At the beginning of an FOPT, cosmic bubbles separating false and true vacua can nucleate stochastically via vacuum decay in field theory. In addition to the well-known channel via quantum tunneling Steinhardt (1982); Coleman (1977); Callan and Coleman (1977), a semiclassical one is also plausible to realize vacuum decay Braden et al. (2019). The former is purely a quantum effect allowing for barrier penetration with lower energy, while the latter could originate from the classical evolution of occasionally developed perturbations Hertzberg and Yamada (2019); Blanco-Pillado et al. (2019); Wang (2019); Huang and Ford (2022). After nucleations, bubbles expand rapidly Cai and Wang (2021); Wang and Yuwen (2023); Wang et al. (2024a); Yuwen et al. (2024) and finally percolate via violent collisions, during which stochastic gravitational wave backgrounds (SGWBs) are generated Caprini et al. (2016, 2020) via bubble collisions Witten (1984); Hogan (1986), sound waves Hogan (1986), and magneto-hydrodynamic (MHD) turbulences Witten (1984) to be detectable by currently running GW detections (LIGO-Virgo Romero et al. (2021); Huang et al. (2021); Jiang and Huang (2023); Badger et al. (2023); Yu and Wang (2023), PTAs Xu et al. (2023); Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023)) and future planning GW detectors (LISA Amaro-Seoane et al. (2017); Baker et al. (2019), Taiji Hu and Wu (2017); Ruan et al. (2020), TianQin Mei et al. (2021); Luo et al. (2016)).
@@
However, the high-frequency tail in the GW power spectrum has not been fully understood yet. Taking the simplest case of vacuum phase transitions without thermal fluids as an example, previous analytical modelings of thin-wall collisions with Jinno and Takimoto (2017) and without Jinno and Takimoto (2019) the envelope approximation agree on an $k^{-1}$ power law for the high-frequency part of the GW spectrum, but a semianalytical simulation from a dubbed bulk flow model Konstandin (2018) also beyond the envelope approximation produces rather different $k^{-2}$ and $k^{-3}$ power laws at ultraviolet (UV) for ultrarelativistic and nonrelativistic walls, respectively. Further relaxing the thin-wall approximation, a previous numerical simulation Cutting et al. (2021); Gould et al. (2021) has found a steeper UV power than $k^{-1}$ depending roughly on a model parameter that is intimately related to the bubble wall thickness.
In this paper, we will make this relation clearer by preliminarily quantifying this relation between the high-frequency power law of GW spectrum and bubble wall thickness at collision. Following Refs. Braden et al. (2015); Gould et al. (2021); Jinno et al. (2019); Lewicki and Vaskonen (2020); Cutting et al. (2021), we first numerically simulate the bubble nucleations, expansion, and collisions in Sec. II and Sec. IV, and then calculate and extract the GW spectrum in Sec. V. In doing so, we analytically derive in Sec. III for the first time the mean bubble separation at nucleation for runaway bubbles from vacuum phase transitions, where bubbles continue to accelerate without approaching a constant terminal velocity at collisions. Regardless of the nucleation channels of quantum or semiclassical types, we find in Sec. VI the high-frequency power law decreases to a certain power of the wall thickness at collisions. Finally in Sec. VII, based on the total energy conservation, we analytically approximate the wall thickness at collisions as a function of other more model-independent parameters, such as the phase hierarchy $\Delta\phi/M_{\mathrm{Pl}}$, strength factor $\alpha=\Delta V/\rho_{\mathrm{rad}}$, inverse duration $\beta/H$, and wall velocity $v_{w}$ as we will define later. This informative modeling of the high-frequency power law together with the remaining parts of the GW spectrum largely breaks model degeneracies for future GW constraints on FOPTs. The last section VIII is devoted to conclusions and discussions.
@@ II Bubble nucleation
In addition to quantum tunneling, bubble nucleation can also occur through a semiclassical mechanism Braden et al. (2019); Hertzberg and Yamada (2019); Blanco-Pillado et al. (2019); Wang (2019); Huang and Ford (2022). In this scenario, quantum fluctuations Hertzberg and Yamada (2019), thermal fluctuations Berera and Fang (1995) or gravitational effects Blanco-Pillado et al. (2019) drive the scalar field or its time derivative to develop a nonzero profile. The classical evolution of these fluctuations then leads to the bubble formation in real space instead of the Euclidean space Braden et al. (2019). One of the simplest realizations Blanco-Pillado et al. (2019); Wang (2019) of this semiclassical mechanism is the time derivative of the scalar field occasionally develops a Gaussian profile large enough to eventually overcome the potential barrier directly from a homogeneous field at the false vacuum. The initial condition for bubble formation is given by
@@
where $A_{0}=\sqrt{2eV_{b}}$ is defined in such a way that the critical velocity $\dot{\phi}_{c}\equiv\sqrt{2V_{b}}$ is realized exactly at $r=R$ for $A=A_{0}$. The profiles of (6) are shown in the right panel of Fig. 2. Ref. Wang (2019) provides a detailed discussion on the ranges of $A/A_{0}$ and $R/R_{0}$ that permit bubble nucleation. In our subsequent calculations, we exactly follow the same configuration and procedure as Ref. Wang (2019) to generate the semiclassical bubble from simply fixing $A/A_{0}=1$ and $R/R_{0}=1$ within the range where bubble nucleation can occur. The specific choice of parameters $A$ and $R$ within this allowable range does not affect our final results.
@@ III Bubble separation
After obtaining the scalar field initial profiles from quantum tunneling and semiclassical mechanisms, there is one last quantity to be determined for the initial configuration of numerical simulation: the mean bubble separation at the nucleation time, which should be twice the mean bubble radius at collisions. In this study, we focus on the vacuum FOPT, and hence, the potential (1) is independent of the background temperature. In such a scenario, the bubble nucleation rate $\Gamma$ is time-independent and given by Coleman (1977); Callan and Coleman (1977); Linde (1981, 1983); Cai et al. (2017b)  [^1] It is important to emphasize that (7) represents only an approximate expression for the bubble nucleation rate, neglecting functional determinant factors and higher-order loop corrections. Specifically, this approximation becomes invalid in the thin-wall limit, for which the correct form of the nucleation rate can be found, for example, in Ref. Garriga (1994). However, since our numerical calculations do not explore the thin-wall limit, the approximation (7) provides a sufficiently accurate description for the scenarios considered in this paper.
$\Gamma\approx\frac{1}{R_{0}^{4}}\left(\frac{S_{E}}{2\pi}\right)^{2}e^{-S_{E}},$ (7)
@@
In the vacuum FOPT, the bubble expands without bound and runs away to approach the speed of light, and hence, there is no constant terminal wall velocity $v_{w}$ by the time of bubble collisions. In this case, the false vacuum fraction $F(t)$ is expressed as Guth and Weinberg (1983); Turner et al. (1992)
@@
where $\Gamma_{\rm E}$ is the Euler-Gamma function and $\Gamma_{\rm E}(8/7)\simeq 0.935$. In Table 1, we present the mean bubble separation for various values of the potential parameter $\omega$. It is observed that, when $\omega<0.6$, the bubble separation becomes extremely large, and simulating such a system would require significant computational resources. Therefore, in this study, we restrict our simulations to values of $\omega$ ranging from 0.6 to 1.0, with an interval of 0.05.
@@
Table 1: The mean bubble distance $d$ at the nucleation time in the quantum phase transition for different $\omega$ values.
Lastly, our analysis above focuses on the mean bubble separation in the quantum phase transition, which is determined by the bubble nucleation rate. For the semiclassical phase transition, the bubble nucleation rate and the corresponding mean separation have been extensively studied in Refs.Braden et al. (2019); Hertzberg et al. (2020); Tranberg and Ungersbäck (2022). In particular, Ref. Braden et al. (2019) has corrected and updated their results in an erratum with a factor of 2 difference in their plotting code as pointed out by Ref. Hertzberg et al. (2020), which causes an enhancement in the nucleation rate by a factor of 1/41/41 / 4 in the slope of $\ln\Gamma$. In this work, when simulating the semiclassical phase transition, we will simply use the same bubble separation values presented in Table 1 but with such a difference in mind that we are not comparing the quantum and semiclassical phase transitions on the same footing.
@@ IV Bubble expansion and collision
but now both evolve in the real space with initial profiles containing two bubbles. For simplicity, we assume the two bubbles are identical. When the bubbles are separated by a distance $d$, their centers can always be positioned on the $z$-axis at $z=\pm d/2$ by adjusting the coordinate system. Now, the evolution of this two-bubble system is best described in cylindrical coordinates $(t,\rho,\theta,z)$, in which case the Eq. (20) is expanded as
@@
[…]
@@
where $\mathrm{d}\mathcal{V}$ is the volume element integrated over the open regions of true bubbles. Throughout the numerical simulations, the total energy $E_{T}\equiv E_{K}+E_{G}+E_{V}$ should remain conserved (or vanished specifically for the quantum phase transition). One of our numerical results for the time evolution of semiclassical bubble expansion is shown in Fig. 3, where it can be seen that $E_{K}$, $E_{G}$, and the absolute value of $E_{V}$ are converged into a constantly evolving total energy $E_{T}$. Specifically, in our simulations, the relative deviation between the total energy $E_{T}$ at the initial and final times is only at the order of $\mathcal{O}(10^{-4})$. For quantum phase transitions, the numerical stability has also been checked with $E_{T}=0$ since the converted potential energy from the false vacuum to true vacuum has all been deposited into the kinetic and gradient energies of the bubble wall.
@@
After the bubble collision, the overlapping region of the two bubbles can rebound back and forth the potential barrier, similar to the trapping effect discussed in Ref. Jinno et al. (2019). When the bubble walls are thick (middle row of Fig. 4), the trapping effect is weakened, but oscillations around the true vacuum become pronounced since the exit point is now on the potential hill away from the true vacuum, generating outgoing waves that follow the expanding bubble walls. As the wall thickness further increases (bottom row of Fig. 4), a significant difference emerges in the semiclassical phase transition, where the scalar field evolution seems to resemble the expansion of a single bubble after collision. This phenomenon occurs as a large wall thickness corresponds to a small bubble separation (as shown in Table 1), causing the classical perturbations to be so close that they can be nearly treated as a single perturbation. In this case, the production of gravitational waves is suppressed, as deviations from sphericity are small, but it is still intriguing for future study as a new way to generate the bubble, not from nucleation, but from the collision of false-vacuum classical perturbations. Nevertheless, in the following Sec. V and Sec. VI, we consider the semiclassical phase transition only for the case where two bubbles remain distinct (corresponding to $\omega<0.9$).
@@ V Gravitational waves
In this work, we consider the system of two identical bubbles, and the cylindrical symmetry in this system allows the integral over the azimuthal angle to be performed analytically. Furthermore, without loss of generality, we can set $\hat{\mathbf{k}}=(\sin\theta_{k},0,\cos\theta_{k})$. The nonzero components of $T_{ij}$ are expressed as
$T_{\rho\rho}(\theta_{k},k)$ $=-\int\mathrm{d}te^{ikt}\int_{0}^{\infty}\mathrm{d}z\cos\left(k\cos\theta_{k}z\right)\int_{0}^{\infty}\mathrm{d}\rho\ \rho\left[\sin^{2}\theta_{k}J_{0}\left(k\sin\theta_{k}\rho\right)+\left(\cos^{2}\theta_{k}+1\right)J_{2}\left(k\sin\theta_{k}\rho\right)\right](\partial_{\rho}\phi)^{2},$
$T_{\rho z}(\theta_{k},k)$ $=-2\int\mathrm{d}te^{ikt}\int_{0}^{\infty}\mathrm{d}z\sin\left(k\cos\theta_{k}z\right)\int_{0}^{\infty}\mathrm{d}\rho\ \rho J_{1}\left(k\sin\theta_{k}\rho\right)\partial_{\rho}\phi\partial_{z}\phi,$
$T_{zz}(\theta_{k},k)$ $=2\int\mathrm{d}te^{ikt}\int_{0}^{\infty}\mathrm{d}z\cos\left(k\cos\theta_{k}z\right)\int_{0}^{\infty}\mathrm{d}\rho\ \rho J_{0}\left(k\sin\theta_{k}\rho\right)(\partial_{z}\phi)^{2},$ (28)
@@
$\omega$ $k_{c}d/2$ $E_{\rm p}$ $A_{m}/A_{n}$ $a_{m}$ $a_{n}$ $l_{\rm col}/R_{\rm col}$ $\dot{l}_{\rm col}$ $v_{\rm col}$ $\sigma_{\rm col}$ 0.60 3.27 1.00 3.32 2.94 1.73 0.0084 0.056 0.974 0.00125 0.65 2.72 0.52 3.12 2.94 1.45 0.0151 0.053 0.968 0.00141 0.70 2.36 0.32 3.12 2.93 1.28 0.0254 0.050 0.956 0.00162 0.75 2.26 0.18 3.21 2.93 1.19 0.0346 0.048 0.948 0.00164 0.80 2.15 0.12 3.32 2.93 1.10 0.0455 0.045 0.941 0.00169 0.85 1.96 0.11 3.23 2.92 1.01 0.0546 0.043 0.933 0.00164
Table 2: The numerical fitting results for the quantum phase transition and semiclassical phase transition. The first column is the potential parameter $\omega$. Parameters from the 2nd to the 6th columns are those shown in the parametrization (30) of the GW spectrum. Here, we set $E_{\rm p}$ in the semiclassical phase transition with $\omega=0.6$ as 1 to normalize the peak value of GW spectra. The 7th column is the ratio of the bubble thickness to its radius at the collision time. The 8th to 10th columns are the wall thickness time derivative, the bubble wall velocity, and the bubble surface tension at the moment of bubble collision.
@@
where $k_{c}$ denotes the peak frequency at which the spectrum reaches its maximum $E_{\rm p}$. Here, $A_{m}$, $A_{n}$, $a_{m}$, and $a_{n}$ characterize the power-law behavior of the GW spectrum. When $k\ll k_{c}$, the spectrum $\Omega_{\rm GW}$ follows the power law $k\mathrm{d}E/\mathrm{d}k\sim(k/k_{c})^{a_{m}}$, whereas for $k\gg k_{c}$, it follows $k\mathrm{d}E/\mathrm{d}k\sim(k/k_{c})^{-a_{n}}$. Numerical fitting results for the parameters in (30) are presented in Table 2. It is easy to see that, at low frequencies, the $a_{m}$ values for different $\omega$ in both phase transitions are nearly identical, $a_{m}\approx 3$, as expected from the causality requirement Cai et al. (2020). At high frequencies, however, the $a_{n}$ values vary depending on $\omega$ and the type of phase transition. Additionally, the peak values $E_{\rm p}$ of the spectrum for both phase transitions decreases as the bubble wall thickness increases, corresponding to higher $\omega$ values.